# Expanding the Resolution Boundary of Outcome-Based Imperfect-Recall Abstraction in Games with Ordered Signals

### v1

@@ Abstract
Keywords: game theory, imperfect-information games, games with ordered signals, computer poker , automated abstraction, hand abstraction, signal abstraction, imperfect-recall
@@ 1 Introduction
Recent years have witnessed significant advancements in AI within the context of Texas Hold’em poker. Pioneering systems like DeepStack (Moravčík et al., 2017), Libratus (Brown and Sandholm, 2018), and Pluribus (Brown and Sandholm, 2019) have demonstrated extraordinary capabilities against professional human players. Much of their success owes to innovative hand abstraction algorithms. These algorithms transform complex hand combinations into simpler, more manageable representative classes, effectively condensing the vast decision space. Such simplification facilitates strategic searches in large-scale games with limited computational resources while maintaining the robustness of the strategies.
Traditionally, Texas Hold’em is modeled as an imperfect-information game, where the task of hand abstraction is categorized as an infoset abstraction task. This modeling paradigm is notably expansive, capturing the essence of Texas Hold’em yet not fully encompassing its nuanced complexities. In Texas Hold’em, hands and actions collectively form the foundational elements that define the infoset, with hand abstraction serving as just one facet of the broader concept of infoset abstraction. The task of hand abstraction poses a unique and independent challenge; however, the model of imperfect-information games fails to capture key concepts such as stages, which are essential in the process of hand abstraction. Previous studies have attempted to engineer the task of hand abstraction from various perspectives, yet this task still necessitates a more precise mathematical model for accurate depiction. Gilpin and Sandholm (2007b) proposed a more detailed game modeling method for this specific type of imperfect-information game, termed ordered signal games. This model skillfully applies the concepts of signals and public action sequences to distinctly outline the two independent components of the infoset in Texas Hold’em-style games: hands and action sequences. However, the model’s reliance on forests to depict the various game stages contributes to its complexity. Moreover, it is confined to describing scenarios where the signal distribution is governed by a combinatorial model with replacement.
@@
In the context of signal abstraction, Gilpin and Sandholm (2007b) developed a lossless abstraction, lossless isomorphism (LI), grounded in the game’s rules and showed that the Nash equilibrium of the abstracted game, derived from LI, can be seamlessly mapped back to the Nash equilibrium of the original game. Although LI offers substantial theoretical insights, its practical implementation in artificial intelligence development is constrained by the large scale of the infosets even after abstraction. Currently, in the domain of hand abstraction for Texas Hold’em, imperfect-recall clustering algorithms based on showdown outcomes have achieved success, with systems like DeepStack, Libratus, and Pluribus employing this method. However, this raises a series of questions:
@@
  • If so, is it possible to expand the resolution boundary of outcome-based imperfect-recall algorithms?
In the second part of this paper, we discuss these questions. Section 6 introduces the concept of potential outcome isomorphism (POI), which aims to construct a signal abstraction based on future showdown outcomes only, striving to identify as many abstracted signal infosets as possible. In Section 7, a theoretical tool called Common Refinement is introduced, which is specifically designed to evaluate the resolution boundaries of signal abstraction algorithms. We demonstrate that POI serves as a common refinement for expected hand strength (E[HS]) and potential-aware methods (PA&PAEMD), the dominant outcome-based imperfect-recall signal abstraction algorithms, thereby effectively establishing the resolution boundary for these methods. Further, we expose a significant resolution gap between POI and LI, highlighting the issue of excessive abstraction present in current outcome-based imperfect-recall algorithms. In Section 8, we analyze the reasons behind the excessive abstraction, pinpointing the neglect of game history sequences as a key factor. To address this, we developed K-Recall Outcome Isomorphism (KROI) by incorporating the historical information, thereby extending the resolution boundary of outcome-based imperfect-recall algorithms. Experimental results indicate that under the one-player-abstraction-perspective, strategies derived from KROI are almost on par with LI and significantly outperform those derived from POI. Even in scenarios of symmetric abstraction, KROI demonstrates substantial strength, significantly surpassing POI. KROI provides us with a new outcome-based imperfect-recall abstraction framework, from which higher-resolution signal abstraction algorithms can be derived.
@@ 1.1 Related Research
Our research is dedicated to hand abstraction techniques in AI systems for Texas Hold’em-style games, a field originating from the works of Shi and Littman (2001) and Billings et al. (2003). These seminal works introduced the concept of game abstraction, aiming to represent games in simplified forms while preserving essential characteristics. Initially, researchers manually categorized hands based on their experience. The first automated approach to hand abstraction was proposed by Gilpin and Sandholm (2006). Subsequently, Gilpin and Sandholm (2007b) introduced games with ordered signals model for Texas Hold’em and developed LI with signal rotation. Despite the elegance of LI, its low compression rates hinder its application in large-scale games, whereas lossy abstraction shows potential for such application. In their work, Gilpin and Sandholm (2007a) proposed expectation-based clustering method and Gilpin et al. (2007) introduced histogram-based clustering method. The former is known as E[HS], while the latter is referred to as potential-aware method. Subsequent studies by Gilpin and Sandholm (2008) and Johanson et al. (2013) compared E[HS] and potential-aware methods, concluding that the latter holds an advantage in large-scale games. Johanson et al. (2013) also introduced the use of earthmover’s distance [^1] https://en.wikipedia.org/wiki/Earth_mover%27s_distance (EMD) in Potential-aware methods. Ganzfried and Sandholm (2014) presented a more efficient approximation algorithm for earthmover’s distance in potential-aware methods, which has been successfully applied in various Texas Hold’em AI systems, marking it as the SOTA work in the field of hand abstraction.
@@
Another area of interest concerns the competitiveness of game-solving strategies after abstraction. Waugh et al. (2009b) noted a non-monotonic relationship between the exploitability of Nash equilibrium in abstracted games and the refinement level of abstraction profiles, a phenomenon termed as abstraction pathology. In a pioneering effort, Sandholm and Singh (2012) developed the first lossy abstraction algorithm that establishes bounds on solution quality for stochastic games. Following this, Kroer and Sandholm (2014) introduced lossy abstraction with quality bounds in general perfect-recall extensive-form games. In practical scenarios, simplified methods with imperfect-recall demonstrate heightened competitiveness, as introduced by Waugh et al. (2009c). Additionally, Lanctot et al. (2012) and Kroer and Sandholm (2016) tailored abstraction algorithms for imperfect-recall extensive-form games under varied conditions, alongside establishing associated solution quality bounds. Essentially, Kroer and Sandholm (2018) unified this collection of research efforts into a cohesive framework while maintaining comparable bounds on solution quality.
@@ 2.1 Texas Hold’em-style Poker Games
Texas Hold’em, a globally popular multiplayer card game, has gained significant prominence in the fields of game theory and artificial intelligence, particularly in Heads-up No-Limit Hold’em (HUNL) variant, which has become a focal point of research. The game of Texas Hold’em unfolds in up to four stages: Preflop, Flop, Turn, and River. At the beginning of each stage, the dealer, representing elements of randomness, draws a specific number of cards from the deck. During the Preflop stage, the dealer distributes two private cards to each player from a standard poker deck, excluding jokers. Subsequently, in the Flop stage, the dealer reveals three community cards from the remaining deck. In the Turn and River stages, one additional community card is revealed in each. Players have the option to raise/bet, call/check, or fold during their turn. If at any point only one player remains, they win all the chips. If multiple players are still present at the end of the River stage without further raises, the pot is allocated based on each player’s betting and the best five-card combination that can be made from his two private cards and the five community cards.
Standard Texas Hold’em, even in Heads-up Limit Hold’em (HULHE) variant, encompasses a vast decision space. To facilitate research, various simplified games have also garnered academic interest, collectively known as Texas Hold’em-style games. For instance, Kuhn Poker (Kuhn, 1950), Leduc Hold’em (Waugh et al., 2009a) and Rhode Island Hold’em (Shi and Littman, 2001) are variants of Texas Hold’em-style games, which employ fewer cards and more streamlined decision-making stages, thereby simplifying the complexity of the game while retaining the core elements. In the experimental section, we introduce Numeral211 Hold’em, a simplified game that closely simulates standard Texas Hold’em with more stages and cards. For details, see Appendix D.
@@ 2.2 Imperfect-Information Games
Texas Hold’em-style games are often described using the model of imperfect-information games, the mathematical specifics of which are detailed in Appendix A. These games involve complex decision-making processes where players may not have complete knowledge of the overall state of the game. In such games, each player, denoted as $i$, makes decisions at non-terminal nodes $h$ in the game, leading to either a terminal node $z$ or another non-terminal node. Each non-terminal node is associated with a specific set of actions $\chi(h)$, and the progression of the game can be influenced by random events controlled by a special player, nature $c$, also called the chance palyer. At terminal nodes $z$, each player receives a specific payoff $u_{i}(z)$. A key concept in imperfect-information games is the infoset $I$, representing the different game nodes indistinguishable to a certain player. Players maintain consistent decision preferences across the nodes within an infoset. The decision preferences of player $i$ across all infosets constitute his strategy $\sigma_{i}$, and the strategies of all players form a strategy profile $\sigma$. The (behavioral) strategy profile influences the probability of reaching each node, whether a non-terminal node $h$ or a terminal node $z$, denoted as $\pi^{\sigma}(h)$ for non-terminal nodes and $\pi^{\sigma}(z)$ for terminal nodes. This, in turn, affects the expected payoff for player $i$ in the game, which is calculated as $v_{i}(\sigma)=\sum_{z\in Z}\pi^{\sigma}(z)u_{i}(z)$. In imperfect-information games, each player aims to maximize his expected payoff.
The solution to an imperfect-information game involves finding a strategy profile $\sigma^{*}$ that is known as a Nash equilibrium (Nash, 1951; Harsanyi, 1995). In a Nash equilibrium, no player can gain additional payoff by unilaterally changing their strategy. Each player’s strategy is the best response to the strategies of others, that is, $v_{i}(\sigma^{*})=\max_{\sigma_{i}\in\Sigma_{i}}v_{i}(\sigma_{i},\sigma^{*}_{-i})$, where $\sigma_{-i}$ represents the strategy profile of all players other than $i$ within the strategy profile $\sigma$. In imperfect-information games, identifying a Nash equilibrium is crucial as it allows players to make optimal decisions amidst uncertainty. To some extent, the Nash equilibrium embodies the optimal solution in these games, especially in the context of two-player scenarios with imperfect-information.
@@
In the field of computational game theory, a widely adopted paradigm for solving large-scale imperfect-information games involves the abstraction of original games into simplified versions by classifying similar states (Sandholm, 2010), as illustrated in Figure 1. This process reduces complexity and facilitates the identification of a Nash equilibrium in the abstracted game, which is then applied back to the original game in hopes of deriving competitive strategies. Typically, the equilibrium found in the abstracted game is expected to closely approximate the Nash equilibrium of the original game. However, this approach has its limitations, one of the main concerns being the abstraction pathology (Waugh et al., 2009b), where the degree of abstraction between the original and the abstracted game does not always correlate monotonically with the distance in their respective Nash equilibrium. This can lead to unpredictable outcomes when applying the equilibrium from the abstracted game to the original one.
Despite these theoretical limitations, this paradigm has proven highly effective in practical applications. For instance, in HUNL, computer programs developed using this approach have surpassed human experts (Brown and Sandholm, 2018, 2019). Hence, this problem-solving paradigm has become a popular and powerful tool in addressing large-scale imperfect-information games.
@@
3 Imperfect-Recall Games and Abstraction
Perfect/imperfect-recall are a pair of characteristics in imperfect-information games. If all players in the game can remember any information they observe during the game, the game is said to have perfect-recall; otherwise, it is said to have imperfect-recall. For a detailed description of this concepts, please refer to Appendix C. Perfect-recall games have several advantages (Kuhn, 1953), with the most important being that the existence of Nash equilibrium is ensured when using behavioral strategies, which are a form of strategy with smaller space cost. However, in practical situations, even if the game does not have perfect-recall, researchers typically use behavioral strategies for strategy solving (Lanctot et al., 2012; Kroer and Sandholm, 2016).
When perfect/imperfect-recall is used to describe abstraction, it refers to whether the abstracted game has perfect/imperfect-recall. However, despite the term imperfect-recall means that players are allowed to forget some memories, in practice, the imperfect-recall abstraction is often used in a more extreme way: requiring players to not retain any memory and make decisions based on future information only (Waugh et al., 2009c; Johanson et al., 2013; Ganzfried and Sandholm, 2014).
@@ 4 Games with Ordered Signals
In imperfect-information games, directly conducting infoset abstraction and action abstraction can be a challenging task. This is because, in the process of infoset abstraction, one must first identify infosets that can be merged, and these infosets must share the same action space. Furthermore, not all infosets with identical action spaces can be combined into the same abstracted infoset. In HUNL, for instance, when the first player in both the Flop and Turn stages faces the same chip pot, their action spaces are identical. However, due to the difference in game stages, it is not suitable to merge these two infosets into the same abstracted infoset.
@@
In practical development of Texas Hold’em AI, we do not simplify the game directly through infoset abstraction and action abstraction. Instead, we consider simplification in two dimensions: hand abstraction and action abstraction. This approach is favored because Texas Hold’em belongs to a special category of imperfect-information games where the description of infosets can be divided into two unrelated dimensions: hands and the action sequences. By focusing on hand abstraction, we can streamline the process of infoset abstraction, allowing us to merge infosets that share the same action sequence. The modeling of this class of games, known as games with ordered signals, was first introduced by Gilpin and Sandholm (2007b).
@@
The original definition of games with ordered signals is somewhat complex. It defines branches at different stages as trees and uses a forest to describe the entire game, involving a significant amount of node mapping. Additionally, its definition of signals is solely based on a combinatorial model without replacement. However, games with ordered signals have the potential for broader applicability. For instance, games similar to Texas Hold’em, such as Liar’s Dice (Freeman, 1989), exhibit significant similarities. Still, due to the distinct nature of signals (with or without replacement), they cannot be modeled using games with ordered signals. In this section, we’ve refined the definition of games with ordered signals for greater clarity and versatility in modeling. We’ve employed this framework to model Texas Hold’em-style games in this paper. Subsequently, our focus in the upcoming sections will be on studying abstraction techniques within this framework.
Definition 1
@@
A game with ordered signals $\tilde{\Gamma}=\left\langle\tilde{\mathcal{N}},\tilde{H},\tilde{Z},\tilde{\rho},\tilde{A},\tilde{\chi},\tilde{\tau},\gamma,\Theta,\varsigma,O,\omega,\succeq,\tilde{u}\right\rangle$ is a structured tuple, where $\tilde{\mathcal{N}}$ is the extended player set, $\tilde{H}$ is the set of non-terminal public nodes, $\tilde{Z}$ is the set of terminal public nodes, $\tilde{\rho}$ is the player function, $\tilde{A}$ is the action set, $\tilde{\chi}$ is the action mapping function, $\tilde{\tau}$ is the successor function, $\gamma$ is the stage function, $\Theta$ is the signal set, $\varsigma$ is the signal revelation function, $O$ is the observation function vector, $\omega$ is the survival status function vector, $\succeq$ is the signal partial order function, and $\tilde{u}$ is the utility function vector. In more detail:
@@
  • $\gamma:\tilde{X}\mapsto\mathbb{N}^{+}$ assigns to each public node $\tilde{x}$ the number of chance public nodes encountered from the initial public node $\tilde{x}^{o}$ to $\tilde{x}$, defined as the stage of $\tilde{x}$. Notably, the stage of every public node is at least 1, indicating that the initial public node is a chance public node.
@@
  • $u=(u_{1},...,u_{N})$, where $u_{i}:\Theta\times\tilde{Z}\mapsto\mathbb{R}$, represents the utility function of player $i\in\mathcal{N}$, which specifies the payoff obtained by player $i$ at terminal public node $\tilde{z}\in\tilde{Z}$ with the signal $\theta\in\Theta$. It is required that for any terminal signal $\theta\in\tilde{\Theta}$ and $\tilde{z}\in\{\tilde{z}^{\prime}\in\tilde{Z}\mid\omega_{i}(\tilde{z}^{\prime})=\omega_{j}(\tilde{z}^{\prime})=true\}$, if $\succeq(\theta,i,j)=true$, then $u_{i}(\theta,\tilde{z})\geq u_{j}(\theta,\tilde{z})$.

### v2

@@ Abstract
keywords:
game theory, imperfect-information games, games with ordered signals, computer poker, automated abstraction, hand abstraction, signal abstraction, imperfect-recall
@@ 1 Introduction
Recent years have witnessed significant advancements in AI within the context of Texas Hold’em poker. Pioneering systems like DeepStack [1], Libratus [2], and Pluribus [3] have demonstrated extraordinary capabilities against professional human players. Much of their success owes to innovative hand abstraction algorithms. These algorithms transform complex hand combinations into simpler, more manageable representative classes, effectively condensing the vast decision space. Such simplification facilitates strategic searches in large-scale games with limited computational resources while maintaining the robustness of the strategies.
Traditionally, Texas Hold’em is modeled as an imperfect-information game, where the task of hand abstraction is categorized as an infoset abstraction task. This modeling paradigm is notably expansive, capturing the essence of Texas Hold’em yet not fully encompassing its nuanced complexities. In Texas Hold’em, hands and actions collectively form the foundational elements that define the infoset, with hand abstraction serving as just one facet of the broader concept of infoset abstraction. The task of hand abstraction poses a unique and independent challenge; however, the model of imperfect-information games fails to capture key concepts such as phases, which are essential in the process of hand abstraction. Previous studies have attempted to engineer the task of hand abstraction from various perspectives, yet this task still necessitates a more precise mathematical model for accurate depiction. Gilpin and Sandholm [4] proposed a more detailed game modeling method for this specific type of imperfect-information game, termed ordered signal games. This model skillfully applies the concepts of signals and public action sequences to distinctly outline the two independent components of the infoset in Texas Hold’em-style games: hands and action sequences. However, the model’s reliance on forests to depict the various game phases contributes to its complexity. Moreover, it is confined to describing scenarios where the signal distribution is governed by a combinatorial model with replacement.
@@
In the context of signal abstraction, Gilpin and Sandholm [4] developed a lossless abstraction, lossless isomorphism (LI), grounded in the game’s rules and showed that the Nash equilibrium of the abstracted game, derived from LI, can be seamlessly mapped back to the Nash equilibrium of the original game. Although LI offers substantial theoretical insights, its practical implementation in artificial intelligence development is constrained by the large scale of the infosets even after abstraction. Currently, in the domain of hand abstraction for Texas Hold’em, imperfect-recall clustering algorithms based on showdown outcomes (outcome-based imperfect-recall abstraction) have achieved success, with systems like DeepStack, Libratus, and Pluribus employing this method. However, this raises a series of questions:
@@
  • If so, is it possible to expand the resolution boundary (which reflects the algorithm’s ability to identify signal information sets) of outcome-based imperfect-recall algorithms?
In the second part of this paper, we discuss these questions. Section 5 introduces the concept of potential outcome isomorphism (POI), which aims to construct a signal abstraction based on future showdown outcomes only, striving to identify as many abstracted signal infosets as possible. In Section 6, a theoretical tool called Common Refinement is introduced, which is specifically designed to evaluate the resolution boundaries of signal abstraction algorithms. We demonstrate that POI serves as a common refinement for expected hand strength (E[HS]) and potential-aware methods (PA&PAEMD), the dominant outcome-based imperfect-recall signal abstraction algorithms, thereby effectively establishing the resolution boundary for these methods. Further, we expose a significant resolution gap between POI and LI, highlighting the issue of excessive abstraction present in current outcome-based imperfect-recall algorithms. In Section 6.3, we analyze the reasons behind the excessive abstraction, pinpointing the neglect of game history sequences as a key factor. To address this, we developed K-Recall Outcome Isomorphism (KROI) by incorporating the historical information, thereby extending the resolution boundary of outcome-based imperfect-recall algorithms. Experimental results indicate that under the one-player-abstraction-perspective, strategies derived from KROI are almost on par with LI and significantly outperform those derived from POI. Even in scenarios of symmetric abstraction, KROI demonstrates substantial strength, significantly surpassing POI. KROI provides us with a new outcome-based imperfect-recall abstraction framework, from which higher-resolution signal abstraction algorithms can be derived.
@@ 1.1 Related Research
Our research is dedicated to hand abstraction techniques in AI systems for Texas Hold’em-style games, a field originating from the works of Shi and Littman [5] and Billings et al. [6]. These seminal works introduced the concept of game abstraction, aiming to represent games in simplified forms while preserving essential characteristics. Initially, researchers manually categorized hands based on their experience. The first automated approach to hand abstraction was proposed by Gilpin and Sandholm [7]. Subsequently, Gilpin and Sandholm [4] introduced games with ordered signals model for Texas Hold’em and developed LI with signal rotation. Despite the elegance of LI, its low compression rates hinder its application in large-scale games, whereas lossy abstraction shows potential for such application. In their work, Gilpin and Sandholm [8] proposed expectation-based clustering method and Gilpin et al. [9] introduced histogram-based clustering method. The former is known as E[HS], while the latter is referred to as potential-aware method. Subsequent studies by Gilpin and Sandholm [10] and Johanson et al. [11] compared E[HS] and potential-aware methods, concluding that the latter holds an advantage in large-scale games. Johanson et al. [11] also introduced the use of earthmover’s distance [^1] https://en.wikipedia.org/wiki/Earth_mover%27s_distance (EMD) in Potential-aware methods. Ganzfried and Sandholm [12] introduced a more efficient approximation algorithm for earthmover’s distance in potential-aware methods (PAEMD). Brown et al. [13] further applied PAEMD to distributed environments for solving large-scale imperfect-information games. This paradigm has found success in Texas Hold’em AI systems and is considered state-of-the-art in hand abstraction.
@@
Another area of interest concerns the competitiveness of game-solving strategies after abstraction. Waugh et al. [14] noted a non-monotonic relationship between the exploitability of Nash equilibrium in abstracted games and the refinement level of abstraction profiles, a phenomenon termed as abstraction pathology. In a pioneering effort, Sandholm and Singh [15] developed the first lossy abstraction algorithm that establishes bounds on solution quality for stochastic games. Following this, Kroer and Sandholm [16] introduced lossy abstraction with quality bounds in general perfect-recall extensive-form games. In practical scenarios, simplified methods with imperfect-recall demonstrate heightened competitiveness, as introduced by Waugh et al. [17]. Additionally, Lanctot et al. [18] and Kroer and Sandholm [19] tailored abstraction algorithms for imperfect-recall extensive-form games under varied conditions, alongside establishing associated solution quality bounds. Essentially, Kroer and Sandholm [20] unified this collection of research efforts into a cohesive framework while maintaining comparable bounds on solution quality.
Other abstraction techniques for decision-making problems include action abstraction [21, 22, 23] and imperfect-recall abstraction [24, 25] in extensive-form games, as well as state abstraction and action abstraction in reinforcement learning [26, 27].
@@ 2.1 Texas Hold’em-style Poker Games
Texas Hold’em is a multiplayer card game that is widespread all over the world. It has gained great prominence in the area of game theory and artificial intelligence, and in turn, the Heads-up No-Limit Hold’em (HUNL) variant it has also gained great prominence and became the main target of the research. Texas Hold’em can be played in up to four phases, namely, Preflop, Flop, Turn, and River. Once the phase begins, the dealer (which corresponds to the random element) deals a certain number of cards from the deck. At Preflop, the dealer deals to every player two private cards from the standard poker deck without jokers. In the next phase, the Flop, the dealer unveals three community cards also from the deck. In the last two phases (Turn and River), another community card is unvealed in each phase. Once it is a player’s turn, he can raise/bet, call/check, or fold. When there is just one player left, he takes all chips. Once the River phase ends with no more raises, the pot is divided according to every player’s betting and the best five-card combination that can be derived from his two private cards and the five community cards.
Standard Texas Hold’em encompasses a huge decision space, even in the Heads-up Limit Hold’em (HULHE) variant. A number of simplified games have been also of academic interest for research purposes and are, all together, called Texas Hold’em-style games: for instance, Kuhn Poker by Kuhn [28], Rhode Island Hold’em by Shi and Littman [5], and Leduc Hold’em by Southey et al. [29], which use fewer cards and more simplified game phases of the decision-making and hence appear less complex but still retain the essential features. In the section of experiments, we introduce a new simplified poker game which we call Numeral211 Hold’em. It closely simulates the standard Texas Hold’em but with more phases and cards than the mentioned ones. More details are available in Appendix D.
@@ 2.2 Imperfect-Information Games
Quite often, Texas Hold’em-style games are described using the model of imperfect-information games, the mathematical specifics of which are detailed in Appendix A. These games are played in complex decision-making processes in which the players possibly have incomplete knowledge of the overall state of the game. In such games, each player, denoted as $i$, makes decisions at non-terminal nodes $h$ in the game, leading to either a terminal node $z$ or another non-terminal node. At each non-terminal node $h$, there exists a set of actions $\chi(h)$ or random events, controlled by a special player, nature $c$, also known as a chance player; at the terminal node $z$, every player receives a specific payoff $u_{i}(z)$. The most important concept in imperfect-information games is the notion of infosets; an infoset $I$ is a set of nodes in the game that are indistinguishable to a certain player, and the player chooses his decision in a consistent manner in every node of an infoset. The decision preferences of player $i$ over all of the infosets is called his strategy $\sigma_{i}$, and of all players - a strategy profile $\sigma$. The (behavioral) strategy profile decides the probability of reaching every node, both non-terminal and terminal, which is denoted as $\pi^{\sigma}(h)$ for non-terminal nodes $h$, and $\pi^{\sigma}(z)$ for terminal nodes $z$. The expected payoff of the player $i$ in the game will depend on the strategy profile and will be: $v_{i}(\sigma)=\sum_{z\in Z}\pi^{\sigma}(z)u_{i}(z)$. In imperfect-information games, every player aims to maximize his expected payoff.
The solution to a game of imperfect information is to find a strategy profile $\sigma^{*}$ that is known as Nash equilibrium [30, 31]. A Nash equilibrium corresponds to the situation in which no player has any incentive, as given, to unilaterally change his own strategy to obtain an extra payoff. Each player’s strategy is a best response to others’ strategies, that is, $v_{i}(\sigma^{*})=\max_{\sigma_{i}\in\Sigma_{i}}v_{i}(\sigma_{i},\sigma^{*}_{-i})$, where $\sigma_{-i}$ is the strategy profile of all players other than $i$ within the strategy profile $\sigma$. In imperfect-information games, identifying a Nash equilibrium is crucial as it allows players to make optimal decisions amidst uncertainty. The Nash equilibrium to some extent resembles the optimal solution in such games, especially when they are two-player with imperfect-information scenarios.
@@
One of the most popular paradigms to follow in the solving of large-scale imperfect-information games in computational game theory is illustrated in Figure 1 [32]. It classifies similar states to abstract original games into more manageable games. This would facilitate the identification of a Nash equilibrium in the abstracted game, which is then applied back to the original game to derive competitive strategies. Generally, the equilibrium found in the abstracted game is expected to be a good approximation of the Nash equilibrium of the original game. However, this approach has some drawbacks, one of the greatest concerns being the abstraction pathology [14], where the degree of abstraction between the original and the abstracted game does not always correlate monotonically with the distance in their respective Nash equilibrium. This can lead to unpredictable outcomes when applying the equilibrium from the abstracted game to the original one.
Despite these theoretical limitations, this paradigm has proven highly effective in practical applications. For instance, in HUNL, computer programs developed using this approach have surpassed human experts [2, 3]. Hence, this problem-solving paradigm has become a popular and powerful tool in addressing large-scale imperfect-information games.
@@
2.4 Imperfect-Recall Games and Abstraction
Perfect/imperfect-recall are a pair of characteristics in imperfect-information games. If all players in the game can remember all information they observed during the game, the game is said to have perfect-recall; otherwise, it is said to have imperfect-recall. For a detailed description of this concepts, please refer to Appendix C. There are several advantages of perfect-recall games [33], the most important of which is that the existence of Nash equilibrium is guaranteed when using behavioral strategies, which is a form of strategy with smaller space cost. However, in practical situations, even if the game does not have perfect-recall, researchers typically use behavioral strategies for strategy solving [18, 19].
The terms perfect/imperfect-recall abstraction are about whether an abstracted game can be made to maintain perfect or imperfect recall. While the term imperfect-recall suggests that players can forget some information, in practice, the main imperfect-recall abstraction method generally refers to an even more extreme concept: players cannot store any memory and may decide based only on future information [17, 11, 12].
@@ 3 Games with Ordered Signals
In imperfect-information games, directly conducting infoset abstraction and action abstraction can be a challenging task. This is because, in the process of infoset abstraction, one must first identify infosets that can be merged, and these infosets must share the same action space. Furthermore, not all infosets with identical action spaces can be combined into the same abstracted infoset. In HUNL, for instance, when the first player in both the Flop and Turn phases faces the same chip pot, their action spaces are identical. However, due to the difference in game phases, it is not suitable to merge these two infosets into the same abstracted infoset.
@@
In practical development of Texas Hold’em AI, we do not simplify the game directly through infoset abstraction and action abstraction. Instead, we consider simplification in two dimensions: hand abstraction and action abstraction. This approach is favored because Texas Hold’em belongs to a special category of imperfect-information games where the description of infosets can be divided into two unrelated dimensions: hands and the action sequences. By focusing on hand abstraction, we can streamline the process of infoset abstraction, allowing us to merge infosets that share the same action sequence. The modeling of this class of games, known as games with ordered signals, was first introduced by Gilpin and Sandholm [4].
@@
The original definition of games with ordered signals is somewhat complex. It defines branches at different phases as trees and uses a forest to describe the entire game, involving a significant amount of node mapping. Additionally, its definition of signals is solely based on a combinatorial model without replacement. However, games with ordered signals have the potential for broader applicability. For instance, games similar to Texas Hold’em, such as Liar’s Dice [34], exhibit significant similarities. Still, due to the distinct nature of signals (with or without replacement), they cannot be modeled using games with ordered signals. In this section, we’ve refined the definition of games with ordered signals for greater clarity and versatility in modeling. We’ve employed this framework to model Texas Hold’em-style games in this paper. Subsequently, our focus in the upcoming sections will be on studying abstraction techniques within this framework.
Definition 1.
@@
A game with ordered signals $\tilde{\Gamma}=\left\langle\tilde{\mathcal{N}},\tilde{H},\tilde{Z},\tilde{\rho},\tilde{A},\tilde{\chi},\tilde{\tau},\gamma,\Theta,\varsigma,O,\omega,\succeq,\tilde{u}\right\rangle$ is a structured tuple, where $\tilde{\mathcal{N}}$ is the extended player set, $\tilde{H}$ is the set of non-terminal public nodes, $\tilde{Z}$ is the set of terminal public nodes, $\tilde{\rho}$ is the player function, $\tilde{A}$ is the action set, $\tilde{\chi}$ is the action mapping function, $\tilde{\tau}$ is the successor function, $\gamma$ is the phase function, $\Theta$ is the signal set, $\varsigma$ is the signal revelation function, $O$ is the observation function vector, $\omega$ is the survival status function vector, $\succeq$ is the signal partial order function, and $\tilde{u}$ is the utility function vector. In more detail:
@@
  • $\gamma:\tilde{X}\mapsto\mathbb{N}^{+}$ assigns to each public node $\tilde{x}$ the number of chance public nodes encountered from the initial public node $\tilde{x}^{o}$ to $\tilde{x}$, defined as the phase of $\tilde{x}$. Notably, the phase of every public node is at least 1, indicating that the initial public node is a chance public node.
@@
  • $\tilde{u}=(\tilde{u}_{1},...,\tilde{u}_{N})$, where $\tilde{u}_{i}:\Theta\times\tilde{Z}\mapsto\mathbb{R}$, represents the utility function of player $i\in\mathcal{N}$, which specifies the payoff obtained by player $i$ at terminal public node $\tilde{z}\in\tilde{Z}$ with the signal $\theta\in\Theta$. It is required that for any terminal signal $\theta\in\tilde{\Theta}$ and $\tilde{z}\in\{\tilde{z}^{\prime}\in\tilde{Z}\mid\omega_{i}(\tilde{z}^{\prime})=\omega_{j}(\tilde{z}^{\prime})=true\}$, if $\succeq(\theta,i,j)=true$, then $\tilde{u}_{i}(\theta,\tilde{z})\geq\tilde{u}_{j}(\theta,\tilde{z})$.